# Stablity of charged scalar hair on a charged black hole

### v1

@@ Abstract
In this paper, we take the Einstein-Maxwell-charged scalar model with an electrically charged scalar field gauge-coupled to the Maxwell field surrounding a charged black hole in the background of a static spherically symmetric metric. We ascertain the validity of all energy conditions coupled with the causality condition, suggesting the possibility of existence of charged hairy solutions. Consequently, we obtain, by exact numerical integration, detailed solutions of the field equations in this model. The solutions exhibit damped oscillatory behaviours for the charged scalar hair. We also find that the electric potential $V(r)$ is a monotonic function of $r$, as required by electrodynamics.
Furthermore, we find that the charged hairy solutions are stable against time-dependant perturbations about the static solutions. We ascertain their stability by carrying out dynamic stability analyses, employing two different methodologies. The first methodology is based on a Sturm-Liouville equation, whereas the second methodology employs a Schrödinger-like equation for the dynamic perturbations. We find stability of the charged hairy solutions in both stability analyses.
@@ 1 Introduction
The Israel-Carter conjecture, or the “no-hair theorem” [1, 2, 3], asserts the uniqueness of the Schwarzschild, Reisnner-Nordström, and Kerr black hole solutions. As a consequence, a black hole can be parametrised only in terms of its mass, electric charge and angular momentum. Early studies in relation to the no-hair theorem in different contexts were made by Bekenstein in the early Seventies [4, 5].
@@
One of the foremost studies in this direction was made by Bekenstein in 1974 [7], which was further extended in Ref. [8]. Bekenstein considered an electrically charged black hole with a massless conformal scalar field non-minimally coupled to gravity via conformal coupling, $\xi R\phi^{2}$. The black hole solution in this model occurs with an event horizon and it is asymptotically flat at infinity. The static spherically symmetric solution was found to be diverging at the horizon. However, upon considering the behaviour of a test particle in this background, it was found that the divergence in the scalar field was not a consequence of an infinite potential barrier and the trajectory of the test particle does not get interrupted at the horizon [8]. Bekenstein’s study established a scalar charge, like other parameters allowed by the no-hair theorem, may also be an admissible parameter in black hole solutions.
Bekenstein’s hairy solution [7, 8] was analysed for stability by Bronnikov and others [9, 10]. Upon analysing a Schröndinger-like equation with an effective potential, it was found that a spherically symmetric perturbation destabilizes the scalar hair. They concluded that the scalar hair decays by evaporation due to this instability, leaving no scalar hair around the black hole.
In an important work, Mayo and Bekenstein [11] considered different combinations of electrically neutral as well as charged black holes with self-interacting neutral and charged scalar fields conformally coupled with gravity via $\xi R|\phi|^{2}$. The charged scalar field was gauge coupled to the Maxwell field. In the case of electrically neutral scalar field, they established a no-hair theorem for the conformal parameter in the ranges $\xi<0$ and $\xi\geq\frac{1}{2}$. They however indicated that the total integrated charge should be unbounded for any value of $\xi$ in the case of charged scalar field.
Bekenstein’s black hole models [7, 8, 11] with conformal scalar hair gave rise to a spate of research studies with modifications. We present below a brief overview of literature in this direction.
Martínez and Zaneli [12] took an electrically neutral black hole with a neutral conformal scalar field in an anti-de Sitter background in $2+1$ dimensions. They obtained a static circularly symmetric hairy black hole solution asymptotically anti-de Sitter at infinity.
Winstanley [13] took a scalar hair model with conformal coupling $\frac{1}{2}\xi R\phi^{2}$ having a potential $V(\phi)=\frac{1}{2}\mu^{2}\phi^{2}$ with no self-interaction, and with a cosmological constant $\Lambda$. Using Bekenstein’s approach [6], existence of hair was ruled out for $\Lambda>0$. However, for $\Lambda<0$, non-trivial hair was found to exist only when $\mu^{2}\leq-\frac{2\Lambda}{3}$.
A generalisation of the above study to dimensions higher than four with both vanishing and non-quadratic scalar potentials $V(\phi)$ was carried out by Radu and Winstanley [14], obtaining stable hairy solutions for $\Lambda<0$. Furthermore, Winstanley [15] examined the conformal model for the massless case ($\mu=0$) and showed that stable hair can exist only when $\Lambda<0$, and for values of conformal parameter in the range $0<\xi\leq\frac{1}{6}$.
Martínez, Troncoso, and Zanelli (MTZ) [16] studied both electrically neutral and charged black holes with a conformally coupled electrically neutral scalar field in $3+1$ dimensions. They showed that hairy solutions can exist for a massless quartically self-interacting scalar field when the cosmological constant is positive. However, Schrödinger like stability analyses of the MTZ solutions, carried out by Harper et al. [17], showed those solutions to be unstable against small scalar perturbations.
Martínez et al. [18] obtained an exact black hole solution with a minimally coupled self-interacting scalar field in an anti-de Sitter background. The external solution has a regular scalar field everywhere. For a given mass, they observed the existence of another branch of solution with vanishing scalar field which could be identified with that in Refs. [19, 20]. This model [18] was further extended to obtain a hairy solution with respect to a parameter occurring in an extended expression for the self-interaction potential of the scalar field in Ref. [21].
Martínez and Troncoso [22] considered an electrically charged black hole with a minimally coupled self-interacting scalar field $\phi$ whose potential has a global maximum at $\phi=0$, the potential behaving like a negative cosmological constant. They obtained a black hole solution with an event horizon determined by the AdS radius and the charge, that gives an upper limit for the charge in terms of the AdS radius. On the other hand, the horizon radius is bounded by the AdS radius, which happens when the charge is zero, in which case the geometry reduces to an AdS spacetime. The scalar hair solution is intimately related with the charge, reducing to a no-hair solution if the charge is zero.
Bardoux et al. [23] studied a charged black hole conformally coupled to a self-interacting electrically neutral scalar field, with a negative cosmological constant. They also coupled the system to a pair of axionic fields that make the event horizon into a planar form, and they showed that the black hole admits a hairy solution. Whereas Martínez et al. [16] showed that a self interacting potential is needed in the presence of a cosmological constant for the existence of a scalar hair at the curved horizon, Bardoux et al. [23] showed that additional axionic fields are required to obtain a hairy solution at the planar horizon.
Campbell et al. [24] took the Kerr black hole geometry minimally coupled to an axion field which gives rise to a source in the Einstein field equation. In addition, the axion potential is sourced by the background geometry. In the slow rotation approximation, they obtained an axionic hairy solution. Campbell et al. [25] further obtained an axion hair solution for a black hole in a low energy string model. Moreover, Lee and Weinberg [26] obtained hairy solution in which an axion like scalar field is non-minimally coupled to electromagnetic field in the gravitational background of a black hole having both electric and magnetic charges.
@@
Poisson [27] examined the black hole solution in Einstein-Gauss-Bonnet gravity in $4+n$ dimensions. Assuming the $n$-dimensional internal space to be slightly perturbed from constant radius, they found that the internal radius behaves like a scalar hair. Subsequent works on scalar hair in Einstein-Gauss-Bonnet gravity can be found in Refs. [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
@@
There also exists a new family of Einstein-Maxwell-scalar field models where an electrically neutral scalar field $\varphi$ is minimally coupled to gravity and non-minimally coupled to the Maxwell field via the interaction $f(\varphi)F^{\mu\nu}F_{\mu\nu}$ [40]. Such models undergo spontaneous scalarization when the effective mass-squared given by $f^{\prime\prime}(0)F^{\mu\nu}F_{\mu\nu}$ changes sign from positive to negative, creating (uncharged) hairy solutions by dynamical growth of the scalar field profile. A comparative study on the choice for $f(\varphi)$ is reported in Ref. [41] and its stability is analysed in Ref. [42]. Extension of such scalarised black hole models with massless [43, 44, 45, 46], massive [47], and massive-self-interacting [48] scalar fields have also been studied.
@@
It is important to mention in this context, and as discussed earlier, that Mayo and Bekenstein [11] had established no-hair theorems in various black hole models with scalar fields conformally coupled to gravity via the interaction $\xi R|\phi|^{2}$. In particular, with the gauge condition of a non-zero electric potential at infinity, they indicated that the total integrated charge should be unbounded. In order to circumvent this situation, they further made the choice of vanishing electric potential $V(r)$ at infinity. However, they found that this latter choice contradicts the fact that $V(r)$ must be a monotonic function of $r$.
In the gauge choice of non-zero electric potential at infinity, it would therefore be important to check whether the electric potential $V(r)$ is a monotonic function of $r$ in such hairy solutions. Furthermore, we must ascertain whether all relevant energy conditions are respected. In addition, we must also ascertain whether such hairy solutions are stable against time-dependant perturbations. This may be relevant in the astrophysical context when a black hole is surrounded by a charged plasma, with its charge electromagnetically coupled to the electric field around the black hole.
@@
In this paper, we therefore consider a Einstein-Maxwell-charged scalar field model where an electrically charged and massive scalar field is gauge-coupled with the Maxwell field giving rise to an interaction of the electric charge $q$ of the scalar field with the electromagnetic field surrounding a charged black hole. We solve this model in a static spherically symmetric spacetime by numerical integration and obtain black hole solutions with electrically charged scalar hair.
We find that the electric potential $V(r)$ is a monotonic function of $r$, as required. We also consider all energy conditions stemming from the energy-momentum tensor as well as the Poynting vector giving the causality condition, and find that all these conditions are satisfied.
@@
We further ascertain the stability of the charged hairy solutions by carrying out dynamic stability analyses employing two different methodologies. The first methodology [49, 50, 51, 52] involves an analysis based on the Sturm-Liouville equation for sinusoidal perturbations about the static hairy solutions. The second methodology [9, 10] involves an analysis of a Schrödinger-like equation for the dynamic perturbations. We find that the charged hairy solutions are stable against dynamical perturbation in both methodologies.
@@ 2 Einstein-Maxwell-charged scalar field model
$\nabla_{\mu}F^{\mu\nu}-2q^{2}A^{\nu}\phi^{*}\phi-q\left(\phi\partial^{\nu}\phi^{*}-\phi^{*}\partial^{\nu}\phi\right)=0.$ (4)
@@
where a prime denotes differentiation with respect to $r$. In addition. from the Einstein field equation 5, we obtain
@@ 3 Energy Conditions for charged Scalar Hair
As we shall see in Sec. 4, we must have the boundary condition $V(r=r_{h})=0$ at the horizon for consistency with the field equations. To analyse the behaviours of the above mixed components at the horizon, we may assume the metric potential to have the Reisnner-Nordström form, $f(r)=1-\frac{2M(r)}{r}+\frac{Q^{2}(r)}{r^{2}}$, and the electric potential $V(r)=\frac{Q(r)}{r}-\frac{Q_{h}}{r_{h}}$. Very close to the horizon, we may write $r=r_{h}+\epsilon$, so that $f(r)=f_{1}\epsilon+O(\epsilon^{2})$ and $V(r)=V_{1}\epsilon+O(\epsilon^{2})$, where $f_{1}=\frac{2M_{h}}{r_{h}^{2}}-\frac{2M_{h}^{\prime}}{r_{h}}-\frac{2Q_{h}^{2}}{r_{h}^{3}}+\frac{2Q_{h}Q_{h}^{\prime}}{r_{h}^{2}}$ and $V_{1}=\frac{Q_{h}^{\prime}}{r_{h}}-\frac{Q_{h}}{r_{h}^{2}}$, giving $\frac{V^{2}(r)}{f(r)}=\frac{V_{1}^{2}}{f_{1}}\epsilon+O(\epsilon^{2})$, which is zero at the horizon ($\epsilon\rightarrow 0$).
Thus all terms in each mixed component given by (17) are finite and well-behaved at the horizon for finite boundary conditions on $\psi(r=r_{h})$, $\psi^{\prime}(r=r_{h})$, together with $\eta(r=r_{h})=0$. Consequently, all energy conditions, supplemented by the causality condition, given by the inequalities in (15) are readily satisfied.
@@ 4 Exact Solutions of the Field Equations
We carry out numerical integrations for the coupled set of nonlinear differential equations (11), (12), (13) and (14) starting from the horizon. In order to set up the boundary conditions for the gradients of the fields, we expand them about the horizon $r=r_{h}$ as
@@
Figure 1: Radial profiles of metric potentials $f(r)$ and $\eta(r)$ with $q=1.0$ and $\mu=0.1$ in units of $r_{h}$. The boundary conditions chosen are $V^{\prime}(r_{h})=1.0$ and $\psi(r_{h})=0.5$. For comparison, the radial profile of the Reisnner-Nordström metric potential $f_{\rm RN}(r)$ is also shown.
@@
The parameter $\epsilon$ is a very small radial displacement from the horizon, which will be controlled in the numerical algorithm. We fix the boundary conditions at the horizon by choosing $V_{1}$, $\psi_{0}$, $q$ and $\mu$, and carry out simultaneous numerical integrations of the coupled set of field equations (11), (12), (13) and (14).
Figure 1 displays the radial profiles of the metric potentials $f(r)$ and $\eta_{(}r)$ with scalar charge $q=1.0$ and mass $\mu=0.1$, in units of $r_{h}$, with the horizon values taken as $V^{\prime}(r_{h})=1.0$, and $\psi_{0}=\psi(r_{h})=0.5$. For comparison with the Reisnner-Nordström solution, we have plotted the profile of $f_{\rm RN}(r)$ with the horizon condition $V^{\prime}(r_{h})=1.0$.
As we can see from Figure 1, the metric coefficient $f(r)$ is slightly non-monotonic near the horizon with a dip at about $r\sim 2r_{h}$, and it acquires a profile similar to $f_{\rm RN}(r)$ for large $r$. There is a suppression in the value of $f(r)$ with respect to $f_{\rm RN}(r)$, which is more pronounced for smaller radial distances.
We also see a steep fall in the metric potential $\eta(r)$ until about $r\sim 2r_{h}$. This correlation originates from the coupled differential equations determining these metric coefficients. Afterwards, $\eta(r)$ appears to approach a constant negative value at infinity, indicating a re-parametrization of the time. These effects are solely due to the presence of the scalar hair. In the absence of scalar hair, $\eta(r)$ would remain zero everywhere.

### v3

# Stability of charged scalar hair on Reissner-Nordström black holes

@@ Abstract
In this paper, we take the Einstein-Maxwell-charged scalar model with an electrically charged scalar field gauge-coupled to the Maxwell field surrounding a charged black hole with a static spherically symmetric metric. In particular, we consider a quadratic scalar potential without any higher order terms and we do not impose any restriction on the magnitude of the scalar charge with respect to the black hole charge.
With this setting, we ascertain the validity of all energy conditions coupled with the causality condition, suggesting the possibility of existence of charged hairy solutions. Consequently, we obtain, by exact numerical integration, detailed solutions of the field equations that incorporate backreaction on the spacetime due to the presence of the charged scalar field. The solutions exhibit damped oscillatory behaviours for the charged scalar hair. We also find that the electric potential is a monotonic function of the radial coordinate, as required by electrodynamics.
In order to ascertain the existence of our charged hairy solutions, we carry out dynamic stability analyses against time-dependant perturbations about the static solutions. For a definite conclusion, we employ two different methodologies. The first methodology involves a Sturm-Liouville equation, whereas the second methodology employs a Schrödinger-like equation, for the dynamic perturbations. We find that our solutions are stable against time-dependant perturbations by both methodologies, confirming the existence of the charged hairy solutions.
@@ 1 Introduction
The Israel-Carter conjecture, or the “no-hair theorem” [1, 2, 3], asserts the uniqueness of the Schwarzschild, Reissner-Nordström, and Kerr black hole solutions. As a consequence, a black hole can be parametrised only in terms of its mass, electric charge and angular momentum. Early studies in relation to the no-hair theorem in different contexts were made by Bekenstein in the early Seventies [4, 5].
@@
One of the foremost studies in this direction was made by Bekenstein in 1974 [7], which was further extended in Ref. [8]. Bekenstein considered an electrically charged black hole with a massless conformal scalar field non-minimally coupled to gravity via conformal coupling, $\xi R\phi^{2}$. This study established a scalar charge, like other parameters allowed by the no-hair theorem, may also be an admissible parameter in black hole solutions. Bekenstein’s hairy solution [7, 8] was analysed for stability by Bronnikov and others [9, 10] and found that a spherically symmetric perturbation destabilizes the scalar hair. They concluded that the scalar hair decays by evaporation due to this instability, leaving no scalar hair around the black hole.
Bekenstein’s black hole models [7, 8] with conformal scalar hair gave rise to a spate of research studies with modifications. Models with conformally coupled scalar fields that are massless [11, 12], massive [13, 14], and self-interacting [15, 16, 17], and with a cosmological constant, were considered to obtain hairy solutions.
@@
Furthermore, Poisson [18] examined the black hole solution in Einstein-Gauss-Bonnet gravity in $4+n$ dimensions. Assuming the $n$-dimensional internal space to be slightly perturbed from constant radius, it was found that the internal radius behaves like a scalar hair. Subsequent works on scalar hair in Einstein-Gauss-Bonnet gravity can be found in Refs. [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Black hole solutions with axionic hair were studied by Campbell et al. [31, 32], Lee and Weinberg [33], and Bardoux et al. [34].
@@
There also exists a new family of Einstein-Maxwell-scalar field models where an electrically neutral scalar field $\varphi$ is minimally coupled to gravity and non-minimally coupled to the Maxwell field via the interaction $f(\varphi)F^{\mu\nu}F_{\mu\nu}$ [35]. Such models undergo spontaneous scalarization when the effective mass-squared given by $f^{\prime\prime}(0)F^{\mu\nu}F_{\mu\nu}$ changes sign from positive to negative, creating (uncharged) hairy solutions by dynamical growth of the scalar field profile. A comparative study on the choice for $f(\varphi)$ is reported in Ref. [36] and its stability is analysed in Ref. [37]. Extension of such scalarised black hole models with massless [38, 39, 40, 41], massive [42], massive-self-interacting [43] scalar fields, and with negative cosmological constant [44] , have also been studied.
@@
In an important work, Mayo and Bekenstein [45] considered different combinations of electrically neutral as well as charged black holes with self-interacting neutral and charged scalar fields conformally coupled with gravity via $\xi R|\phi|^{2}$. The charged scalar field was gauge coupled to the Maxwell field. In the case of electrically neutral scalar field, they established a no-hair theorem for the conformal parameter in the ranges $\xi<0$ and $\xi\geq\frac{1}{2}$. They however indicated that the total integrated charge should be unbounded for any value of $\xi$ in the case of charged scalar field. In order to circumvent this situation, they further made the choice of vanishing electric potential $V(r)$ at infinity. However, they found that this latter choice contradicts the fact that $V(r)$ must be a monotonic function of $r$.
The no scalar hair theorem by Mayo and Bekenstein [45] was further examined for charged scalar fields gauge coupled with Maxwell field by Hong et al. [46, 47]. It was pointed out that the contribution of the scalar mass in the asymptotic solution was not properly taken into consideration in the analysis of Mayo and Bekenstein. They considered a charged scalar field gauge coupled to the Maxwell field with polynomial interacting potential with scalar charge being very small compared to the charge of the black hole. Such a system with a charged scalar cloud around the black hole was studied without any backreaction to the Reissner-Nordström background in Ref. [46] and with backreaction in Ref. [47]. The latter model was further studied by Herdeiro and Radu [48] for Schwarzschild and Reissner-Nordström black holes with charged clouds as hair. They examined various conditions for the existence of charged scalar hair.
A common feature in these models was to take the scalar potential to be a polynomial of sixth degree and the scalar charge to be much smaller than the black hole charge. Consequently, we may ask the question whether a charged scalar hair would exist as a stable solution if the scalar field is taken to be a charged massive Klein-Gordon field (with quadratic potential) without any limitation to its charge compared to the black hole charge.
@@
In this paper, we therefore consider a Einstein-Maxwell-charged scalar field model where an electrically charged and massive Klein-Gordon field (with quadratic potential) is gauge-coupled to the Maxwell field giving rise to an interaction of the electric charge $q$ of the scalar field with the electromagnetic field surrounding a charged black hole. Importantly, we take account of the backreaction on the exterior spacetime due to the presence of the scalar field by considering all field equations without making any approximation. We solve this model in a static spherically symmetric spacetime by numerical integration and obtain black hole solutions with electrically charged scalar hair. We also consider all energy conditions stemming from the energy-momentum tensor as well as the Poynting vector giving the causality condition, and find that all these conditions are satisfied.
@@
We further ascertain the stability of the charged hairy solutions by carrying out dynamic stability analyses employing two different methodologies. The first methodology [49, 50, 51, 52] involves an analysis based on the Sturm-Liouville equation for sinusoidal perturbations about the static hairy solutions. The second methodology [9, 10] involves an analysis of a Schrödinger-like equation for the dynamic perturbations. We find that the charged hairy solutions are stable against dynamical perturbations in both methodologies.
@@ 2 Einstein-Maxwell-charged scalar field model
$\nabla_{\mu}F^{\mu\nu}-2q^{2}A^{\nu}\phi^{*}\phi-iq\left(\phi\partial^{\nu}\phi^{*}-\phi^{*}\partial^{\nu}\phi\right)=0.$ (4)
@@
where a prime denotes differentiation with respect to $r$. Here $V^{\prime}(r)$ is the magnitude of the electric field due to electric charges of both the black hole as well as the scalar field enclosed by a sphere of radius $r$. In addition, from the Einstein field equation (5), we obtain
@@ 3 Energy conditions for charged scalar hair
Figure 1: Radial profiles of the mixed components of energy-momentum tensor $|T^{t}_{\ t}(r)|$, $|T^{r}_{\ r}(r)|$, $|T^{\theta}_{\ \theta}(r)|$, and $|T^{\varphi}_{\ \varphi}(r)|$ with $q=1.0$ and $\mu=0.1$ in units of $r_{h}$. The boundary conditions chosen are $V^{\prime}(r_{h})=1.0$ and $\psi(r_{h})=0.5$.
@@
As we shall see in Sec. 4, we must have the boundary condition $V(r=r_{h})=0$ at the horizon for consistency with the field equations. To analyse the behaviours of the above mixed components at the horizon, we may assume the metric potential to have the Reissner-Nordström form, $f(r)=1-\frac{2M(r)}{r}+\frac{Q^{2}(r)}{r^{2}}$, and the electric potential $V(r)=\frac{Q(r)}{r}-\frac{Q_{h}}{r_{h}}$. Very close to the horizon, we may write $r=r_{h}+\epsilon$, so that $f(r)=f_{1}\epsilon+O(\epsilon^{2})$ and $V(r)=V_{1}\epsilon+O(\epsilon^{2})$, where $f_{1}=\frac{2M_{h}}{r_{h}^{2}}-\frac{2M_{h}^{\prime}}{r_{h}}-\frac{2Q_{h}^{2}}{r_{h}^{3}}+\frac{2Q_{h}Q_{h}^{\prime}}{r_{h}^{2}}$ and $V_{1}=\frac{Q_{h}^{\prime}}{r_{h}}-\frac{Q_{h}}{r_{h}^{2}}$, giving $\frac{V^{2}(r)}{f(r)}=\frac{V_{1}^{2}}{f_{1}}\epsilon+O(\epsilon^{2})$, which is zero at the horizon ($\epsilon\rightarrow 0$).
Thus all terms in each mixed component given by (17) are finite and well-behaved at the horizon for finite boundary conditions on $\psi(r=r_{h})$, $\psi^{\prime}(r=r_{h})$, together with $\eta(r=r_{h})=0$. Consequently, all energy conditions, supplemented by the causality condition, given by the inequalities in (15) are readily satisfied. In order to illustrate the validity of these energy conditions, we show in Figure 1 the radial profiles of the mixed components of the energy momentum tensor.
@@ 4 Exact solutions of the field equations
We carry out numerical integrations for the coupled set of nonlinear differential equations (11), (12), (13) and (14) starting from the horizon.
4.1 Horizon boundary conditions
In order to set up the boundary conditions for the gradients of the fields, we expand them about the horizon, $r=r_{h}$, as
@@ 4.1 Horizon boundary conditions
The parameter $\epsilon$ is a very small radial displacement from the horizon, which will be controlled in the numerical algorithm.
On the other hand, we expect a Minkoskian spacetime as $r\rightarrow\infty$ so that $f\rightarrow 1$. Moreover, with the gauge choice of $V=0$ at the horizon, $V(r)$ and $\eta(r)$ approach constant values as $r\rightarrow\infty$. With these asymptotic behaviours, equation (14) reduces to $\psi^{\prime\prime}+\frac{2}{r}\psi^{\prime}+\omega^{2}\psi=0$ at infinity, where $\omega^{2}=q^{2}e^{2\eta_{\infty}}V_{\infty}^{2}-\mu^{2}=$ constant. The corresponding solutions that satisfy the above asymptotic requirements is $\psi\propto\frac{\cos(\omega r)}{r}$ for $\omega^{2}>0$ and $\psi\propto\frac{e^{-\omega r}}{r}$ for $\omega^{2}<0$. Both solutions vanish as $r\rightarrow\infty$. However, the latter solution is not possible in the present model with boundary conditions discussed above.
4.2 Numerical integration
Employing the boundary conditions at the horizon as discussed in the previous subsection, we carry out simultaneous numerical integrations of the coupled set of field equations (11), (12), (13) and (14) by choosing the parameters $V_{1}$, $\psi_{0}$, $q$ and $\mu$, so as to obtain exact solutions for the radial profiles of $\psi(r)$, $V(r)$, $f(r)$, and $\eta(r)$, until the radial distance from the black hole horizon is sufficiently high, where the scalar field approaches a vanishing value. The results of the numerical integration are presented below.
Figure 2: Radial profiles of metric potentials $f(r)$ and $\eta(r)$ with $q=1.0$ and $\mu=0.1$ in units of $r_{h}$. The boundary conditions chosen are $V^{\prime}(r_{h})=1.0$ and $\psi(r_{h})=0.5$. For comparison, the radial profile of the Reissner-Nordström metric potential $f_{\rm RN}(r)$ is also shown.
Figure 2 displays the radial profiles of the metric potentials $f(r)$ and $\eta_{(}r)$ with scalar charge $q=1.0$ and mass $\mu=0.1$, in units of $r_{h}$, with the horizon values taken as $V^{\prime}(r_{h})=1.0$, and $\psi_{0}=\psi(r_{h})=0.5$. For comparison with the Reissner-Nordström solution, we have plotted the profile of $f_{\rm RN}(r)$ with the horizon condition $V^{\prime}(r_{h})=1.0$.
Our numerical data indicates that $f_{\rm RN}=0.95$ and $f=0.86$ at $\frac{r}{r_{\rm h}}=30$, $f_{\rm RN}=0.98$ and $f=0.94$ at $\frac{r}{r_{\rm h}}=100$, and $f_{\rm RN}=0.99$ and $f=0.96$ at $\frac{r}{r_{\rm h}}=200$. Thus the metric coefficient $f(r)$ approaches the limiting value of unity slowly (compared to the Reissner-Nordström case) due to backreaction of the charged scalar hair. It is however clear that the flat spacetime limit would be reached upon carrying out the numerical computation up to sufficiently high values of the radial coordinate $r$.
As we can see from Figure 2, the metric coefficient $f(r)$ is slightly non-monotonic near the horizon with a dip at about $r\sim 2r_{h}$, and it acquires a profile similar to $f_{\rm RN}(r)$ for large $r$. We also see a steep fall in the metric potential $\eta(r)$ until about $r\sim 2r_{h}$. This correlation originates from the coupled differential equations determining these metric coefficients. Afterwards, $\eta(r)$ appears to approach a constant negative value at infinity, indicating a re-parametrization of the time. These effects are solely due to the presence of the scalar hair. In the absence of scalar hair, $\eta(r)$ would remain zero everywhere.
Figure 3 shows the radial profiles of the scalar field $\psi(r)$ for different values of mass $\mu$ and charge $q$ of the scalar field, with $V^{\prime}(r_{h})=1.0$ and $\psi(r_{h})=0.5$ at the horizon. All these profiles show a damped oscillatory behaviour with respect to the radial distance $r$. The profile with $\mu=0.1$ and $q=1.0$ shows a steep fall until $r\sim 2r_{h}$. We note that this steep fall in $\psi(r)$ is correlated with the non-monotonic behaviour of $f(r)$ and a steep fall of $\eta(r)$ as noted previously in Figure 2.